A $10,000 corporate bond has a 4% stated interest rate making the semiannual interest payments $200 for each half year period. The bond has eight years to maturity. If current market rates for interest are 3% per year, what should you pay for this bond?
Answer #1

face value = 10000

Semi annual payment = 200

Market rate = 3% per year = 3%/2 = 1.5% per semi annual period

Maturity = 8 years = 8*2 = 16 semi annual periods

Present value of the bond = 200 * (P/A, 1.5%, 16) + 10000*(P/F,1.5%, 16)

= 200 * 14.131264 + 10000 * 0.788031

= 10706.56